OpenStudy (anonymous):

I'm doing an Algebra II project and I don't understand one of the steps... Can someone help me?

OpenStudy (anonymous):

what is it on?

OpenStudy (anonymous):

V=lwh l=4in w=3in h=in My specific formula is 12=4*3*1. Here's the step I don't get: Rewrite the formula using the variable x for the length. Substitute the value of the volume found in step 2 for V and express the width and height of the object in terms of x plus or minus a constant. For example, if the height measurement is 4 inches longer than the length, then the expression for the height will be (x + 4).

OpenStudy (anonymous):

Would I have 12=(x-1)(x-3)?

OpenStudy (anonymous):

@amistre64 @ash2326 @Hero @Mandre

OpenStudy (anonymous):

@emilyjm the topic is polynomials

hero (hero):

For volume of rectangular prism, you need to multiply length, width, and height V = lwh Of course, you know that.. Now, if you want to create an algebraic equation based on it, you need to define two of the variables in terms of the other variable. Say, you wanted length to be x, you could say that the width is 2 units greater than the length (x + 2)...the height is 4 units less than the length (x - 4) and the Volume is 84 so: 84 = x(x + 2)(x - 4)

OpenStudy (anonymous):

okay...

hero (hero):

either way, there needs to be three specific multiplications on the right side all in terms of one variable.

hero (hero):

It is good that you are trying to do this project on your own.

OpenStudy (anonymous):

@Hero so I would write it as 12=x(x-1)(x-3)

hero (hero):

Should be...

OpenStudy (anonymous):

Thank you sooo much!

OpenStudy (anonymous):

@Hero how would I write this in standard form? That's the next step... I thought standard form was ax^2+bx+c=0... When I simplify it, I get x^3...